The assets (in billions of dollars) of the four wealthiest people in a particular country are 35, 34, 28, 17. Assume that samples of size n 2 are randomly selected with replacement from this population of four values a. After identifying the 16 different possible samples and finding the mean of each sample, construct a table representing the sampling distribution of the sample mean In the table, values of the sample mean that are the same have been combined Probability Probability 35 34.5 34 31.5 31 28 26 25.5 22.5 17 Type integers or fractions.) b. Compare the mean of the population to the mean of the sampling distribution of the sample mean The mean of the population, (Round to two decimal places as needed.) c. Do the sample means target the value of the population mean? In general, do sample means make good estimates of population means? Why or why not? The sample means Vthe mean of the sample means ?| the population mean. In general, sample means | ?I make good estimates of population means because the mean is V estimator
Answer #1

for number of samples =4*4=16

hence proabbiltiy of each sample =1/16

xbar probability xbar probability
35    1/16 28    1/16
34.5    1/8 26    1/8
34    1/16 25.5    1/8
31.5    1/8 22.5    1/8
31    1/8 17    1/16

b)

mean of the population 28.5 is equal to the mean of the sample means 28.5

c)

the sample means target the population mean. in genereal sample means do make good.........the mean is unbiased estimator.